OpenStudy (anonymous):

write a quadratic function that fits the points (0,5), (3,8), & (4,13). << correct answer : f(x)=x^2-2x+5 >>

OpenStudy (ranga):

The general form of q quadratic equation is: y = ax^2 + bx + c You have three unknowns here: a, b and c. You are given 3 points. Put the x,y values of each point in the equation and you will have 3 equations. Three unknowns, three equations. You can solve for a, b and c.

OpenStudy (ranga):

First point is (0, 5). Put x = 0, y = 5 in y= ax^2 + bx + c 5 = 0 + 0 + c So c = 5 So our quadratic equation now is: ax^2 + bx + 5 Second point is (3,8). Put x = 3, y = 8 in y = ax^2 + bx + 5 8 = a(9) + 3b + 5 9a + 3b = 3 -------- equation (1) Third point is (4, 13). Put x = 4, y = 13 in y = ax^2 + bx + 5 13 = a(16) + 4b + 5 16a + 4b = 8 -------- equation (2) Eliminate b from equations 1 and 2 Multiplying equation 1 by 4: 36a + 12b = 12 -------- equation (3) Multiplying equation 2 by 3: 48a + 12b = 24 -------- equation (4) Subtract equation 3 from 4: 12a = 12 So a = 1 Put a=1 in equation 3: 36 + 12b = 12 12b = -24 b = -2 So we have a = 1, b = -2 and c = 5 for the equation y = ax^2 + bx + c y = (1)x^2 + (-2)x + (5) y = x^2 - 2x + 5

OpenStudy (anonymous):

oh my god thank you so much

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

Is this Algebra 2 or linear algebra?

OpenStudy (anonymous):

Alg II

OpenStudy (anonymous):

okay good to know. @ranga did a good job.

OpenStudy (ranga):

Thanks @wio